# Why Transformers Need Adam: A Hessian Perspective

### v1

@@ 2.5 Implication on Choosing SGD or Adam
The above findings could bring up an interesting empirical guidance: we can compute the blockwise spectrum of initial Hessian, and then decide whether to use Adam or SGD. Such a method could be useful in scenarios in training large models that are not mainstream Transformers or CNNs, e.g., Mamba (Gu and Dao, 2023). In these cases, there is not much prior experience in choosing optimizers. It would be intriguing to decide whether SGD is suitable for the task before the training is launched. We present two reasons.
  • In large-scale model training, memory is limited but the model size grows, so Adam could become a major computational burden due to the memory of 2nd-order momentum (as mentioned in Section 1 and (Rasley et al., 2020)).
  • SGD is a lightweight alternative to Adam, but the performance of SGD may be bad. In large NN training, it might take days (or even weeks) to obtain informative feedback from training logs, so it is expensive to conduct trial and error to choose an appropriate optimizer. Specifically, if SGD performs badly, it would be difficult to judge whether the bad performance is due to improper learning rates or other more fundamental reasons. Then, it would be hard to decide whether to switch the optimizer.
Motivated by the above reasons, we here propose a quantitative metric that could predict the failure of SGD before the training is launched. With the help of this metric, we could save much expense on the trial and error for SGD. The metric is simply the averaged JS distance among blockwise Hessian spectra at initialization, i.e., the averaged value in each heatmap of Figure LABEL:fig_blockwise_heatmap. We denote it as $JS^{0}$. We present $JS^{0}$ of various models in Table 1. We note that $JS^{0}$ establishes a quantitative difference between the loss landscape of Transformers and CNNs. As argued in the previous sections, this quantity can predict the failure of SGD. Note that $JS^{0}$ is independent of optimizers and could be efficiently checked in priori before training.
Model ResNet18 VGG16 ViT-base BERT GPT2-nano GPT2
$JS^{0}$ 0.10 0.09 286.41 53.38 14.93 83.23
Table 1: $JS^{0}$ denotes the average JS distance between the initial Hessian spectra of each pair of parameter blocks. We find that the $JS^{0}$ on CNNs is at least 140×140$\times$140 × smaller than on Transformers. $JS^{0}$ establishes a quantitative difference between the loss landscape of Transformers and CNNs.
To summarize this section, we provide an explanation for SGD’s failure. We identify a phenomenon called block heterogeneity. This phenomenon naturally exists in Transformers, but not in CNNs. This phenomenon can exist in non-Transformer architectures as well. Our experiments suggest a strong connection between the failure of SGD and block heterogeneity. Our findings could provide empirical guidance to decide whether to use Adam or SGD.
@@ 3 Case Study of Quadratic Models and Preliminary Theory
Now we construct quadratic functions with block diagonal Hessian. We will compare the performance of GD and Adam on Hessian with and without block heterogeneity. Initial theoretical results on these quadratic models will be provided. Note that insights on quadratic models could be important for understanding realistic NNs, as mentioned by researchers such as LeCun et al. (2002) and OpenAI team (Kaplan et al., 2020).
3.1 Settings and Additional Notations
We consider $\min_{w}\mathcal{L}(w)=\frac{1}{2}w^{T}Hw-h^{T}w$ where $H\in\mathbb{R}^{d\times d}$ is positive definite and $h\in\mathbb{R}^{d}$. We denote $\mathcal{L}^{*}$ as the minimum value of $\mathcal{L}(w)$. We set $H$ as a block diagonal matrix: $H=\operatorname{diag}(H_{1},\cdots,H_{L})$, where $H_{l}\in\mathbb{R}^{d_{l}\times d_{l}}$ and $d=\sum_{l=1}^{L}d_{l}$. We use $w_{l}\in\mathbb{R}^{d_{l}}$ to denote the variable in the $l$-th block and $w=(w_{1}^{T},\cdots,w_{L}^{T})^{T}\in\mathbb{R}^{d}$. Similarly for $h_{l}\in\mathbb{R}^{d_{l}}$. Similarly, we use $[\nabla L(w)]_{l}\in\mathbb{R}^{d_{l}}$ to denote the gradient in the $l$-th block and denote $[\mathcal{L}(w)]_{l}=\frac{1}{2}(w_{l}^{t})^{T}H_{l}w_{l}^{t}-h_{l}^{T}w_{l}$ as the objective function w.r.t. the $l$-th block. Note that $\mathcal{L}(w)=\sum_{l=1}^{L}[\mathcal{L}(w)]_{l}$. We denote $\lambda_{1}\geq\lambda_{2}\cdots\geq\lambda_{d}$ as the eigenvalues of $H$. Similarly for $\lambda_{l,1}\cdots\lambda_{l,d_{l}}$. We denote $\kappa=\frac{\lambda_{1}}{\lambda_{d}}$ and $\kappa_{l}=\frac{\lambda_{l,1}}{\lambda_{l,d_{l}}}$ as the condition number of $H$ and $H_{l}$, respectively. We say an algorithm has complexity $\tilde{\mathcal{O}}(C)$ if it takes $\mathcal{O}(C\log(1/\epsilon))$ iterations to achieve error $\frac{\mathcal{L}(w)-\mathcal{L}^{*}}{\mathcal{L}\left(w^{0}\right)-\mathcal{L}^{*}}\leq\epsilon$, where $w^{0}$ is the initial point.
@@ 3.2 Experimental Observations
For all experiments, we choose $h=0$. We consider four types of Hessian $H$ as follows.
  • Case 1: Hessian with Transformer-type spectra. We choose $L=4$ and $d_{l}=100$. For $l\in[L]$, we construct $H_{l}=Q_{l}\Lambda_{l}Q_{l}^{T}$ where $Q_{l}$ are independent standard Gassian random matrix and $\Lambda_{l}$ are diagonal matrices. For the diagonal elements in $\Lambda_{l}$, we sample $d_{l}$ numbers according to the spectrum of the embedding layer; 3rd Query, 3rd Value, 3rd MLP (fc layer) in GPT2. Shifting and proportional scaling are performed to ensure all elements in $\Lambda_{l}$ lie in the interval $[1,2000]$. This ensures strong convexity and controls the condition number of $H$ equals $2000$. The spectra of $H_{l}$ are in Figure LABEL:fig:heter-block.
  • Case 2: Hessian with CNN-type spectra. We consider the same setup as Case 1. For the diagonal elements in $\Lambda_{l}$, we sample $d_{l}$ numbers according to the spectrum of the 1st to 4th convolution layers in ResNet18. We shift and proportionally scale the elements $\Lambda_{l}$ to the interval $[1,2000]$ to ensure the strong convexity and the condition number $H$ equals 2000. The spectra of $H_{l}$ are shown in Figure LABEL:fig:homo-block.
  • Case 3: Hessian with simplified heterogeneous spectra. We choose $L=3$ and $d_{l}=3$. For $l\in[L]$, we construct $H_{l}=Q_{l}\Lambda_{l}Q_{l}^{T}$ where $Q_{l}$ are independent standard Gassian random matrix and $\Lambda_{l}$ are diagonal matrices. We set the diagonal elements of $\Lambda_{l}$ as $\{1,2,3\},\{99,100,101\},\{1998,1999,2000\}$ for $l=1,2,3$, respectively. The spectra of $H_{l}$ are different due to their different supports. The condition number of Hessian $H$ is $2000$.
  • Case 4: Hessian with simplified homogeneous spectra. We consider the same setup as Case 3. We set the diagonal elements of $\Lambda_{l}$ as $\{1,99,1998\},\{2,100,1999\},\{3,101,2000\}$ for $l=1,2,3$, respectively. The spectra of $H_{l}$ are similar. The condition number of Hessian $H$ is 2000.
Note that in all four cases, the Hessian share the same condition number. We study two types of optimizers: one that assigns a single learning rate for all blocks, and ones that assign different learning rates across blocks. The optimizers are introduced as follows.
@@
    We use the optimal learning rate $\eta=\frac{2}{\mu+L}$ (Nesterov, 2013). We use a standard Gaussian random vector as initialization.
  • Coordinate-wise-learning-rate optimizer. We study Adam with a constant learning rate and with no bias correction for simplicity (Algorithm 3). We set $\beta_{1}=0$ to erase the effect of momentum. This helps us to focus on the effect of coordinate-wise learning rate (or the effect of diagonal preconditioning) in Adam. We use $\epsilon=0$. We consider $\beta_{2}=1$ and $\beta_{2}=0.999$, respectively. When $\beta_{2}=1$, Adam assigns coordinate-wise learning rates according to the initial gradient, but these learning rates are fixed along iteration. The update rule is as follows.
@@
    where $D_{Adam}^{0}=\operatorname{diag}(\nabla\mathcal{L}(w^{0})\circ\nabla\mathcal{L}(w^{0}))^{\frac{1}{2}}$ and $\nabla\mathcal{L}(w^{0})=Hw^{0}-h$. When $\beta_{2}<1$, the coordinate-wise learning rates adaptively change along iteration. The update rule is as follows.
    $w^{t+1}=w^{t}-\eta(D_{Adam}^{t})^{-1}\nabla\mathcal{L}(w)=w^{t}-\eta(D_{Adam}^{t})^{-1}(Hw^{t}-h),$ (3)
    where $D_{Adam}^{t}=\operatorname{diag}\left((1-\beta_{2})\left(\sum_{k=1}^{t}\beta_{2}^{t-k}\nabla\mathcal{L}(w^{k})\circ\nabla\mathcal{L}(w^{k})\right)+\beta^{t}\operatorname{diag}(\nabla\mathcal{L}(w^{0})\circ\nabla\mathcal{L}(w^{0}))\right)^{\frac{1}{2}}$ and $\nabla\mathcal{L}(w^{k})=Hw^{k}-h$. We grid search $\eta$ and use a standard Gaussian random vector as initialization.
    We remark that when $\beta_{2}<1$, Adam would repeatedly bounce among non-optimal points, causing non-convergence. This will be shown in Proposition 2.
Summary of experimental observations.
We now compare the performance of Adam and GD on these four types of Hessians. The results are shown in Figure LABEL:fig_quadratic. For Hessian with heterogeneous blocks like Case 1 and 3, GD largely lags behind Adam. For Hessian with homogeneous blocks (Case 2 and 4), GD is on par or even better than Adam. We emphasize that all Hessians have the same condition number in these four cases. Further, Hessian of Case 3 and 4 share all the eigenvalues (not just the extreme ones). The performance gap between Adam and GD is due to the different blockwise spectra caused by the different locations of eigenvalues. We hypothesize that GD performs badly because it uses one single learning rate for all blocks, which cannot handle the heterogeneity among blocks. Such heterogeneity can be better handled using different learning rates across blocks, as designed in Adam.
@@ 3.3 Initial Theoretical Results
Unfortunately, the blockwise spectrum is rarely discussed in the optimization area. A most related notion is “block Lipschitz constant" (Beck and Tetruashvili, 2013) for studying block coordinate descent (BCD) type methods, but it was not linked to the performance of SGD or Adam before. To our knowledge, we are not aware of any theory of Adam or GD built on the block diagonal structures or the blockwise spectra of Hessian. Further, there seems to be no result distinguishing the complexity of single-learning-rate methods and the coordinate-wise counterpart, either. We now make an initial attempt in this direction. We first present the lower bound for GD.
@@
Proposition 1 shows that GD has complexity $\tilde{\mathcal{O}}(\kappa)$ and such complexity is tight. Now we prove that: if we are allowed to choose a customized learning rate for each block sub-matrix $H_{l}$, then the complexity will be improved. We call this method blockwise GD (BGD). Note that BGD is impractical to implement since it is hard to choose $L$ learning rates in advance. BGD serves as an artificial method for theoretical understanding.
Blockwise GD (BGD).
Choose learning rates $\{\eta_{l}\}_{l=1}^{L}$ for all blocks in advance and performs the following:
$w^{t+1}=w^{t}-(D_{BGD})^{-1}\nabla\mathcal{L}(w)=w^{t}-(D_{BGD})^{-1}(Hw^{t}-h),$
where $D_{BGD}=\operatorname{diag}\left(\left(\underbrace{\eta_{1},\cdots,\eta_{1}}_{d_{1}\text{times}},\cdots,\underbrace{\eta_{L},\cdots,\eta_{L}}_{d_{L}\text{times}}\right)^{T}\right)\in\mathbb{R}^{d\times d}.$
BGD can be also viewed as a diagonally preconditioned version of GD. We now show that this diagonal preconditioner can significantly improve the complexity of GD.
@@
(Upper bound for GD with blockwise learning rate.) Consider $\min_{w}\mathcal{L}(w)=\frac{1}{2}w^{T}Hw-h^{T}w$ where $H\in\mathbb{R}^{d\times d}$ is positive definite and $h\in\mathbb{R}^{d}$. We assume $H$ is a block diagonal matrix, i.e., $H=\operatorname{diag}(H_{1},\cdots,H_{L})$, where $H_{l}\in\mathbb{R}^{d_{l}\times d_{l}}$ and $d=\sum_{l=1}^{L}d_{l}$. Let $w_{BGD}^{t}$ be the output of BGD after $t$ steps. There exists a set of $\{\eta_{l}\}_{l=1}^{L}$ such that:
$\mathcal{L}(w_{BGD}^{t+1})-\mathcal{L}^{*}\leq\max_{l\in[L]}\left(1-\frac{1}{\kappa_{l}}\right)\left(\mathcal{L}(w_{BGD}^{t})-\mathcal{L}^{*}\right)$ (5)
where $\kappa_{l}$ is the condition number of $H_{l}$.
Theorem 1 shows that: if we are allowed to choose a learning rate for each $H_{l}$, the complexity could be $\tilde{O}(\max_{l\in[L]}\kappa_{l})$, in lieu of $\tilde{O}(\kappa)$ as in GD. That is to say, the rate of BGD only depends on the "slowest block". Note that in certain heterogeneous block Hessian such Case 1 and ${\bf 3}$, $\kappa_{l}$ is about 20×20$\times$20 × and 2000×2000$\times$2000 × smaller than $\kappa$, respectively. In these cases, blockwise-learning-rate methods can be about 20×20$\times$20 × and 2000×2000$\times$2000 × faster than single-learning-rate methods like GD, respectively.
Despite the advantages of BGD, it is impractical to find the proper learning rates for all blocks, especially when the number of blocks $L$ is large. Fortunately, these learning rates could be approximated by the preconditioner in Adam, e.g., $D_{Adam}^{0}$ in (2). We now provide a complexity bound of Adam.
Assumption 1.
(Bounded initialization.) The initialization is bounded and is far from the optimal solution:
$C_{l,1}\lambda_{l,1}\leq\left|[\nabla\mathcal{L}(w^{0})]_{l,i}\right|\leq C_{l,2}\lambda_{l,1},\text{ for }i\in[d_{l}],l\in[L].$
Theorem 2.
(Upper bound for Adam with $\beta_{2}=1$.) Consider the same setting as in Theorem 1 and consider Adam with $\beta_{1}=0$ and $\beta_{2}=1$ as in (2). Assume the initialization satisfies Assumption 1. Let $w_{Adam}^{t}$ be the output of Adam after $t$ steps. Let $\eta=\min_{l\in[L]}\frac{1}{C_{l,1}}$. We have
@@
where $\kappa_{Adam,l}=r\kappa_{l}$, $r:=\max_{l\in[L]}\left(\frac{C_{l,2}}{C_{l,1}}\right)^{2}$ and $\kappa_{l}$ is the condition number of $H_{l}$.
The proofs of all the above theorems are shown in Appendix E. Theorem 2 states that Adam (with $\beta_{2}=1$) has complexity $\tilde{O}\left(r\cdot\max_{l\in[L]}\kappa_{l}\right)$. We remark that condition $\beta_{2}=1$ is necessary because any $\beta_{2}<1$ causes non-convergence issue (Da Silva and Gazeau, 2020; Bock and Weiß, 2019). We restate their results in Proposition 2. The non-convergence is also observed in Figure LABEL:fig_quadratic (c), where we find that the iterates of Adam quickly converge to near-optimal solutions, and then bounce back. As such, $\beta_{2}=1$ is necessary for asymptotic analysis. As shown in (Da Silva and Gazeau, 2020), the non-convergence is due to the constant learning rate.
@@
(Non-convergence of constant-learning-rate Adam with $\beta_{2}<1$.) (Da Silva and Gazeau, 2020, Proposition 12) Consider $\min_{w\in\mathbb{R}}\mathcal{L}(w)=\frac{1}{2}w^{2}$. Consider Adam with $\beta_{1}=0$ and $\beta_{2}<1$ as in (3). Let $w_{Adam}^{t}$ be the output of Adam after $t$ steps. There exists a discrete limit cycle for (3), i.e., Adam produces oscillations and ${\small\liminf_{t\rightarrow\infty}\left(\mathcal{L}(w_{Adam}^{t})-\mathcal{L}^{*}\right)>0}$.
We now compare the complexity of Adam and that of GD. By Theorem 2, Adam is faster than GD when $r\cdot\max_{l\in[L]}\kappa_{l}\leq\kappa$. In the quadratic model with heterogeneous blocks (Case 3), our simulation over 1000 trials shows that $r\leq 100$ with probability $\geq\frac{2}{3}$ when using standard Gaussian random initialization. Since $\max_{l\in[L]}\kappa_{l}\approx 1$, we have $r\cdot\max_{l\in[L]}\kappa_{l}\leq 100$, w.h.p., and is 20×20$\times$20 × smaller than $\kappa=2000$. So Adam could be 20×20$\times$20 × faster than GD, w.h.p.. This is indeed observed in Figure LABEL:fig_quadratic where Adam outperforms GD by a significant margin. We summarize the complexity of GD, BGD, and Adam in Table 2.
Optimizer GD BGD Adam with Adam with

### v2

@@ 2.6 Implication on Choosing SGD or Adam
The above findings show that SGD can perform badly on various architectures. This leads to an intriguing question: Can we predict the bad performance of SGD before training begins?
Our findings can bring up an empirical guidance: we can compute the blockwise spectrum of initial Hessian, and then decide whether to use Adam or SGD. Such a method could be useful in scenarios in training large models that are not mainstream Transformers or CNNs, e.g., Mamba [38]. In these cases, there is not much prior experience in choosing optimizers. It would be intriguing to decide whether SGD is suitable for the task before the training is launched. One might argue that simple trial is enough: try both SGD and Adam; if Adam is remarkably better, then pick Adam; if Adam and SGD are similar, then pick SGD. Nevertheless, this simple approach may not be easy for large models. First, for large models, it may take days to know one run of an algorthm is good or not. Second, it requires tuning hyperparameters at least a few times to get a reasonably good judgement, making the cost of trial even higher.
We here propose a quantitative metric that could predict the incompetence of SGD before the training. With the help of this metric, we could save much expense on the trial and error for SGD. The metric is simply the averaged JS distance among blockwise Hessian spectra at initialization, i.e., the averaged value in the heatmap of Figure LABEL:fig_blockwise_heatmap. We denote it as $JS^{0}$. We present $JS^{0}$ of various models in Table 1. Note that $JS^{0}$ establishes a quantitative difference between the loss landscape of Transformers and CNNs. Further, $JS^{0}$ is independent of optimizers and could be checked before training.
Table 1: $JS^{0}$ denotes the average JS distance between the initial Hessian spectra of each pair of parameter blocks. A larger $JS^{0}$ suggests that the task is more difficult for SGD.
Model ResNet18 VGG16 GPT2 (pretrained) MLP-mixer BERT GPT2 ViT-base
$JS^{0}$ 0.10 0.09 18.84 34.90 53.38 83.23 286.41
To validate the effectiveness of the quantitative metric $JS^{0}$, we summarize $JS^{0}$ of different models and the corresponding SGD performance in Figure 4. We find that the performance gap between SGD and Adam becomes greater as $JS^{0}$ increases. Thus, $JS^{0}$ can serve as a potential indicator to predict whether SGD may underperform Adam.
Figure 4: Comparison of $JS^{0}$ and the performance of SGD on different models. We find the performance gap between SGD and Adam becomes greater as $JS^{0}$ increases.
@@ 3 Case Study of Quadratic Models and Preliminary Theory
Now we study quadratic functions with block diagonal Hessian, with or without block heterogeneity. Note that insights on quadratic models could be important for understanding realistic NNs, as mentioned by researchers such as LeCun et al. [49] and OpenAI team [44].
Setups and additional notations. We consider the following quadratic minimization.
$\min_{w\in\mathbb{R}^{d}}\mathcal{L}(w)=\frac{1}{2}w^{T}Hw-h^{T}w,$
where $H\in\mathbb{R}^{d\times d}$ is positive definite and $h\in\mathbb{R}^{d}$. We denote $\mathcal{L}^{*}$ as the minimum value of $\mathcal{L}(w)$. We set $H$ as a block diagonal matrix: $H=\operatorname{diag}(H_{1},\cdots,H_{L})$, where $H_{l}\in\mathbb{R}^{d_{l}\times d_{l}}$ and $d=\sum_{l=1}^{L}d_{l}$. We use $w_{l}\in\mathbb{R}^{d_{l}}$ to denote the variable in the $l$-th block and $w=(w_{1}^{T},\cdots,w_{L}^{T})^{T}\in\mathbb{R}^{d}$. Similarly for $h_{l}\in\mathbb{R}^{d_{l}}$. Similarly, we use $[\nabla L(w)]_{l}\in\mathbb{R}^{d_{l}}$ to denote the gradient in the $l$-th block and denote $[\mathcal{L}(w)]_{l}=\frac{1}{2}(w_{l}^{t})^{T}H_{l}w_{l}^{t}-h_{l}^{T}w_{l}$ as the objective function w.r.t. the $l$-th block. Note that $\mathcal{L}(w)=\sum_{l=1}^{L}[\mathcal{L}(w)]_{l}$. We denote $\lambda_{1}\geq\lambda_{2}\cdots\geq\lambda_{d}$ as the eigenvalues of $H$. Similarly for $\lambda_{l,1}\cdots\lambda_{l,d_{l}}$. We denote $\kappa=\frac{\lambda_{1}}{\lambda_{d}}$ and $\kappa_{l}=\frac{\lambda_{l,1}}{\lambda_{l,d_{l}}}$ as the condition number of $H$ and $H_{l}$, respectively. We say an algorithm has complexity $\tilde{\mathcal{O}}(C)$ if it takes $\mathcal{O}(C\log(1/\epsilon))$ iterations to achieve error $\frac{\mathcal{L}(w)-\mathcal{L}^{*}}{\mathcal{L}\left(w^{0}\right)-\mathcal{L}^{*}}\leq\epsilon$, where $w^{0}$ is the initial point.
@@ 3.1 Experimental Observations
We consider four types of Hessian $H$ as follows. For all cases, we set condition number = 5000.
  • Case 1: Hessian with Transformer-type spectra. We choose $L=4$ and $d_{l}=25$. For $l\in[L]$, we construct $H_{l}=Q_{l}\Lambda_{l}Q_{l}^{T}$ where $Q_{l}$ are matrices with i.i.d. standard Gassian entries and $\Lambda_{l}$ are diagonal matrices. For the diagonal elements in $\Lambda_{l}$, we sample $d_{l}$ numbers according to the spectrum of the embedding layer; 3rd Query, 3rd Value, 3rd MLP (fc layer) in GPT2. Shifting and proportional scaling are performed to ensure all elements in $\Lambda_{l}$ lie in the interval $[1,5000]$. This ensures strong convexity and controls the condition number of $H$ equals $5000$. The spectra of $H_{l}$ are in Figure LABEL:fig:heter-block in Appendix B. We choose $h=0$ for all cases.
  • Case 2: Hessian with CNN-type spectra. We use the same setup as in Case 1. For the diagonal elements in $\Lambda_{l}$, we sample $d_{l}$ numbers according to the spectrum of the 1st to 4th convolution layers in ResNet18. We then shift and scale $\Lambda_{l}$ to the interval $[1,5000]$ to ensure strong convexity and a condition number of 5000. The spectra of $H_{l}$ are shown in Figure LABEL:fig:homo-block in Appendix B.
  • Case 3: Hessian with simplified heterogeneous spectra. We choose $L=3$ and $d_{l}=3$. For $l\in[L]$, we construct $H_{l}=Q_{l}\Lambda_{l}Q_{l}^{T}$ where $Q_{l}$ are independent standard Gassian random matrix and $\Lambda_{l}$ are diagonal matrices. We set the diagonal elements of $\Lambda_{l}$ as $\{1,2,3\},\{99,100,101\},\{4998,4999,5000\}$ for $l=1,2,3$, respectively. The spectra of $H_{l}$ are different due to their different supports. The condition number of Hessian $H$ is $5000$.
  • Case 4: Hessian with simplified homogeneous spectra. We consider the same setup as Case 3. We set the diagonal elements of $\Lambda_{l}$ as $\{1,99,4998\},\{2,100,4999\},\{3,101,5000\}$ for $l=1,2,3$, respectively. The spectra of $H_{l}$ are similar. The condition number is 5000.
Now we study two types of optimizers: one that assigns a single learning rate for all blocks, and one that assign different learning rates across blocks.
@@
    We use the optimal learning rate $\eta=\frac{2}{\mu+L}$ [63]. We use standard Gaussian initialization.
  • Coordinate-wise-learning-rate optimizer. We study Adam with a constant learning rate and with no bias correction for simplicity (Algorithm 3). We set $\beta_{1}=0$ to erase the effect of momentum. This helps us to focus on the effect of coordinate-wise learning rate (or the effect of diagonal preconditioning) in Adam. We use $\epsilon=0$. We consider $\beta_{2}=1$ and $\beta_{2}=0.99$, respectively. When $\beta_{2}=1$, Adam assigns coordinate-wise learning rates according to the initial gradient, but these learning rates are fixed along iteration. The update rule is as follows.
@@
    where $D_{Adam}^{0}=\operatorname{diag}(\nabla\mathcal{L}(w^{0})\circ\nabla\mathcal{L}(w^{0}))^{\frac{1}{2}}$ and $\nabla\mathcal{L}(w^{0})=Hw^{0}-h$. When $\beta_{2}<1$, the coordinate-wise learning rates adaptively change along iteration. The update rule is as follows (note that $\nabla\mathcal{L}(w^{k})=Hw^{k}-h$.).
    $w^{t+1}=w^{t}-\eta(D_{Adam}^{t})^{-1}\nabla\mathcal{L}(w)=w^{t}-\eta(D_{Adam}^{t})^{-1}(Hw^{t}-h),\quad\text{where}$ (3)
    $D_{Adam}^{t}=\operatorname{diag}\left((1-\beta_{2})\left(\sum_{k=1}^{t}\beta_{2}^{t-k}\nabla\mathcal{L}(w^{k})\circ\nabla\mathcal{L}(w^{k})\right)+\beta^{t}\operatorname{diag}(\nabla\mathcal{L}(w^{0})\circ\nabla\mathcal{L}(w^{0}))\right)^{\frac{1}{2}}$
    We grid search $\eta$ and use the standard Gaussian initialization. We remark that when $\beta_{2}<1$, Adam would bounce among non-optimal points. This will be shown in Proposition 2.
Summary of experimental observations. Figure LABEL:fig_quadratic presents two phenomena. For Hessian with heterogeneous blocks (Case 1 and 3), GD largely lags behind Adam. For Hessian with homogeneous blocks (Case 2 and 4), GD is on par with Adam. We emphasize that all Hessians have the same condition number. Further, Hessian of Case 3 and 4 share all the eigenvalues (not just the extreme ones). The gap between Adam and GD is due to the different blockwise spectra caused by the different locations of eigenvalues. We hypothesize that GD performs badly because it uses one single learning rate for all blocks, which cannot handle the heterogeneity among blocks. Such heterogeneity can be better handled using different learning rates across blocks, as designed in Adam.
@@ 3.2 Initial Theoretical Results
Note that the blockwise spectrum is not common in the optimization area. The most related notion is perhaps “block Lipschitz constant" [8] for studying block coordinate descent (BCD) type methods, but it was not linked to the performance of SGD or Adam before. To our knowledge, we are not aware of any theory of Adam or GD built on the block diagonal structures or the blockwise spectra of Hessian. We now make an initial attempt in this direction. We first present the lower bound for GD.
@@
Proposition 1 shows that GD has complexity $\tilde{\mathcal{O}}(\kappa)$ and such complexity is tight. Now we prove that Adam can achieves better complexity. This is because it chooses different learning rates for different block sub-matrix $H_{l}$ via its diagonal preconditinoner $D_{Adam}^{0}$. We consider generic random initialization that covers commonly used distributions such as Gaussian, Uniform, etc.
Assumption 1.
(Random initialization.) Assume the initialization $w^{0}$ is sampled from a continuous distribution, i.e., the probability measure (induced by $w^{0}$) of any zero-Lebesgue-measure set is 0.
@@
(Upper bound for Adam with $\beta_{2}=1$.) Consider the same setting as Proposition 1 and consider Adam with $\beta_{1}=0$ and $\beta_{2}=1$ as in (2). Assume the initialization satisfies Assumption 1. Let $w_{Adam}^{t}$ be the output of Adam after $t$ steps. Let $\eta=\min_{l\in[L]}\frac{1}{C_{l,1}}$. Then w.p.1., we have
@@
where $\kappa_{Adam,l}=r\kappa_{l}$, $\kappa_{l}$ is the condition number of $H_{l}$, constant $r$ relates to $w^{0}$ defined as:
$r=\frac{\max_{l\in[L]}C_{l,2}^{2}}{\min_{l\in[L]}C_{l,1}^{2}},\text{ where }C_{l,1}=\min_{i\in[d_{l}]}\frac{|[\nabla\mathcal{L}(w^{0})]_{l,i}|}{\lambda_{l,1}},C_{l,2}=\max_{i\in[d_{l}]}\frac{|[\nabla\mathcal{L}(w^{0})]_{l,i}|}{\lambda_{l,1}}.$ (6)
The proofs of the above theorems are shown in Appendix E. Theorem 1 states that Adam (with $\beta_{2}=1$) has complexity $\tilde{O}\left(r\cdot\max_{l\in[L]}\kappa_{l}\right)$. We note that coefficient $r$ depends on the ratio between initial gradient and the principal eigenvalue for each block, and smaller ratio would give faster convergence. We further remark that condition $\beta_{2}=1$ is necessary because any $\beta_{2}<1$ causes non-convergence issue [10, 21]. We restate their results in Proposition 2. The non-convergence is also observed in Figure LABEL:fig_quadratic (c), where we find that the iterates of Adam quickly converge to near-optimal solutions, and then bounce back. As such, $\beta_{2}=1$ is necessary for asymptotic analysis. The analysis for $\beta_{2}=1$ is still meaningful since it still shows the effect of Adam’s preconditioner.
As shown in [21], the non-convergence is due to the constant learning rate. Reducing the learning rate reduces the gap between $\mathcal{L}(w_{Adam}^{t})$ and $\mathcal{L}^{*}$, but does not remove it.
@@
(Non-convergence of constant-learning-rate Adam with $\beta_{2}<1$.) [21, Proposition 12, Figure 1] Consider $\min_{w\in\mathbb{R}}\mathcal{L}(w)=\frac{1}{2}w^{2}$. Consider Adam with $\beta_{1}=0$ and $\beta_{2}<1$ as in (3). Let $w_{Adam}^{t}$ be the output of Adam after $t$ steps. There exists a discrete limit cycle for (3) and $\liminf_{t\rightarrow\infty}\left(\mathcal{L}(w_{Adam}^{t})-\mathcal{L}^{*}\right)>0$.
We now compare the complexity of Adam and that of GD. By Theorem 1, Adam is faster than GD when $r\cdot\max_{l\in[L]}\kappa_{l}\leq\kappa$. In the quadratic model with heterogeneous blocks (Case 3), our simulation over 1000 trials shows that $r\leq 1000$ with probability $\geq\frac{2}{3}$ when using standard Gaussian random initialization. Since $\max_{l\in[L]}\kappa_{l}\approx 1$, we have $r\cdot\max_{l\in[L]}\kappa_{l}\leq 1000$, w.h.p., and is about 5×5$\times$5 × smaller than $\kappa=5000$. So Adam could be 5×5$\times$5 × faster than GD, w.h.p.. This is indeed observed in Figure LABEL:fig_quadratic where Adam outperforms GD by a significant margin. We summarize the complexity of GD and Adam in Table 2. We remark that there is room to improve the complexity bound of Adam, but it could be challenging. We provide more technical discussions in Appendix B.
Table 2: The complexity of GD and Adam for minimizing a strongly convex quadratic function with block diagonal Hessian. The symbol ✗ means non-convergence. $\kappa$ and $\kappa_{l}$ denote the condition number of the full Hessian and the block submatrix, respectively. $r$ is defined in (6).
Optimizer GD Adam with Adam with